# Dynamical complex networks for understanding motility induced phase separation in Active Brownian Particles

### v1

@@ 4.2 Transient Region: Hybrid phase
We define the transient region as the intermediate phase between the one-phase and two-phase regions (blue-green in Fig. 2). Since it is an intermediate phase, it is expected to have a degree of distribution with a hybrid behavior. Physically and analyzing our numerical simulations, we note that in this regime, small clusters composed of three disks (triangular arrangement) appear and eventually break because of the persistent ABP motion. In Fig. 6, we observe the degree distribution, average path length ($l_{G}$), and average clustering coefficient ($\bar{C}$) at times $t=0.45t_{r}$ and $t=11t_{r}$. Interestingly, the initially random graph (Gaussian distribution) evolves to a hybrid distribution with two components (Gaussian and power law). This hybrid behavior tells us that the system is trying to evolve from a random to a scale-free graph, but the values for the packing fraction and Péclet are not large enough. Moreover, at $t=11t_{r}$, the system does not reach the steady state, which is not the case for the one-phase and two-phase regions. The latter tells us that this hybrid phase is more complex to analyze regarding its geometrical and topological properties since the system remains out of equilibrium.
Figure 6: Top: Degree distribution for the case at $\text{Pe},\phi=80,0.4$ at the initial time $t=0.45t_{r}$ (right) and time $t=11.0t_{r}$ (left). The inset shows the same data plotted in $\log{}-\log{}$. Bottom: Average path length (left ) and average clustering coefficient (left) over time of the system. Both show an increasing behavior over time until $\approx 8t_{r}$.
4.3 Two-phase region: Phase separation
@@
At more significant packing fractions and Péclet numbers, i.e., $\phi\geq 0.5$ and $\mbox{Pe}\geq 80$, the system enters the inner region of Fig. 2, where MIPS is manifested. From the point of view of the ABP, as time evolves, the initially random distribution (liquid-like state) evolves to a state where a gas-like and a solid-like phase cohabit in the system (green-red regions in Fig. 2). In terms of the complex network, the degree distribution of the system evolves from a Gaussian distribution at short times ($t=0.45t_{r}$) towards a double power law distribution for larger times ($t\geq 11t_{r}$); see Fig. 7(a)(b). Notably, when the degree distribution is plotted in the log-log scale, two slopes adequately describe the distribution $P(k)$, see Fig. 7(c)(d). This double power law distribution has the form
@@
where $K=6$, $k_{c}\in[13-17]$ is the critical degree and $(\gamma_{1},\gamma_{2})$ are the critical exponents. The value $K=6$ is connected with the 2D disk packing problem, where each disk has exactly six nearest neighbors in an ideal cluster and corresponds to the peak of $P(k)$, see Fig. 7(b). Moreover, the critical values for $k_{c}$ depend on the clustering coefficient and Péclet number. The first slope $\gamma_{1}$ (red dashed line of Fig. 7) changes from values close to $1.5$ at $\phi=0.5$ towards values close to $3.7$ at $\phi=0.7$. The second slope $\gamma_{2}$ (black dashed lines) varies from values close to $12$ to $17$. In summary, we obtain $\gamma_{1}\approx(1.5-3.7)$ and $\gamma_{2}\approx(12-17)$, and therefore $\gamma_{1}<\gamma_{2}$ in our ABP model. In Table 1, we compare our results with different systems that exhibit a double power law distribution.
@@
The temporal behavior of both curves is shown in Fig. 8 ($\gamma_{1}$) and ($\gamma_{2}$). We note that $\gamma_{1}$ is more stable over time, and the packing fraction dominates its steady value in some inverse relation. On the contrary, $\gamma_{2}$ shows more oscillations, which can be explained because the clustering formation is unstable for the values of $(\phi,\mbox{Pe})$.
@@
Figure 7: Degree distribution for phase-separated systems. (a) $\log{}-\log{}$ plot shows a clear Gaussian distributions at time $t=0.45t_{r}$. (b) degree distribution at $t=11.0t_{r}$ shows an evolution in the distribution where the degree $k=6$ for each curve is the one with the highest probability. Figures (c) and (d) show two distributions of the figure (b) zoomed in at $\text{Pe},\phi=(80,0.7)$ and $\text{Pe},\phi=(140,0.5)$ respectively, where the red and black dashed lines are the best linear fit for the first and second power law decay respectively for each distribution.
Figure 8: Slopes of the power law decay in the degree distribution for phase-separated systems. The initial data point occurs when the degree $k=6$ has the highest probability, indicating distinct curves for each system. Consequently, we observe different transition times. Notably, around $t\approx 8t_{r}$, the value of $\gamma_{1}$ (top) stabilizes consistently across all systems. In contrast, the values for the second slope, $\gamma_{2}$ (bottom), appear noisy and do not seem to stabilize.
4.3.1 Role of the average clustering coefficient
We can characterize this phase separation using the network average clustering coefficient $\bar{C}$ given in Eq. 9. In Fig. 9, we plot $\bar{C}$ as a function of time for different values of the packing fraction and Péclet number for the frontier and inner regions of the phase diagram illustrated in Fig. 2. To better understand the observed values of $\bar{C}$, we can elaborate a simple theoretical analysis. The clustering coefficient can be computed as $C=\lambda_{G}(v)/\tau_{G}(v)$, where $\lambda_{G}(v)$ is the number of triangles with one vertex on $v$ and $\tau_{G}(v)$ is the number of triples with one vertex on $v$ ($v$ is a vertex of the graph $G$). For a simple hexagonal cluster composed of seven particles (see inset of Fig. 10 (left)), we can calculate the clustering coefficient for one particle inside the cluster ($C_{i}$) and one particle in the border ($C_{b}$). Using this reasoning, we obtain $C_{i}=6/15=0.4$ and $C_{b}=2/3\approx 0.6667$. Let be $\bar{C}_{hex,N_{r}}$ the average clustering coefficient for a hexagonal arrangement with $N_{r}$ concentric hexagonal rings. For $N_{r}$ hexagonal rings, we have $6N_{r}$ particles in the border and $1+3N_{r}(N_{r}+1)$ particles in total. Therefore, the fraction of particles in the border and inside the cluster are $f_{b}(N_{r})=6N_{r}/(1+3N_{r}(N_{r}+1))$ and $f_{i}=(1+3N_{r}(N_{r}-1))/(1+3N_{r}(N_{r}+1))$, respectively. Thus, we can define the average clustering coefficient as,
$\bar{C}_{hex,N_{r}}=f_{b}(N_{r})C_{b}+f_{i}(N_{r})C_{i}={2\over 15}\left[5-2f_{i}(N_{r})\right],$ (14)
where we have used $f_{b}+f_{i}=1$. For instance, for $N_{r}=1$ (seven particles) we get $\bar{C}_{hex,7}=(1/7)C_{m}+(6/7)C_{b}\approx 0.6286$ (first value of Fig. 10 (left)). However, as one increases the fractions of particles inside the cluster ($f_{i}\rightarrow 1$), we get $\bar{C}_{hex,N_{r}}\rightarrow C_{i}=0.4$ (asymptotic value of Fig. 10 (left)). Also, it is important to mention that three disks forming a minimal cluster have an average cluster coefficient $\bar{C}=1$ (maximum value).
In Fig. 9, we note that in the frontier region of the MIPS phase diagram ($\phi=\{0.3,0.4\}$ and Pe =$=\{80,140\}$), where small triangular clusters randomly appear, we observe $0.2\lesssim\bar{C}\lesssim 0.35$, and we associate these values for very low probability events (probabilities $\sim(0.2-0.35)$) of forming small triangular clusters. However, as we enter the inner region of the MIPS phase diagram ($\phi=\{0.5,0.6,0.7\}$ and Pe =$=\{80.140\}$), finite hexagonal clusters are dominant, explaining the observed values $0.35\lesssim\bar{C}\lesssim 0.42$ at large times oscillates around $0.4$. A steady state value $\bar{C}\approx 0.35$ can be explained by the fact that a perfect hexagonal cluster with a clustering coefficient of 0.4 can be penalized by the fraction of particles with respect to the total number of particles. A value $\bar{C}\approx 0.42$ can be understood as the reference value 0.4 with a small contribution of small triangular clusters.
Figure 9: Top: Network average clustering coefficient for systems in the frontier region (top) and the inner region (bottom). The values for $\bar{C}$ for one-phased systems(top) at $\phi=0.3$ have virtually no change over time and slowly increase at $\phi=0.4$. In contrast, $\bar{C}$ for the two-phased systems increases rapidly at first and then stabilizes around $t=8t_{r}$, showing a tendency to $\bar{C}=0.4$.
Figure 10: Left: Network average clustering coefficient of identical circles in a hexagonal packing arrangement, using $C_{r}=2^{1/6}\sigma$ ( black circles). $\bar{C}$ approaches asymptotically to $0.4$ (dashed red line), which is the value of the clustering coefficient of the center particle in the seven-particle arrangement (inset). Right: Average path length for the same hexagonal packing arrangement as a function of the packing fraction $\phi$ (black circles). We see that $l_{G}$ grows as $30\sqrt{\phi}$ (red dashed line) and could represent an upper bound with respect to $\phi$ for clustered ABP systems.
4.3.2 Role of the average path length
@@
The average path length defined in Eq. 10 was measured for both regions, the frontier and inner region. In Fig. 11, we observe the time evolution for the network average path length in the case of a system with a single phase characterized by a Gaussian degree distribution $l_{G}$ has a constant value for different Pe numbers, meanwhile on the transient region for $\phi=0.4$ and Pe=$80$ the value increases slowly not reaching a constant value see blue dots in Figure 11.
In Fig. 10, we can observe a particular scaling for the average path length. In the ideal case (perfect hexagonal clusters with $N$ particles), we note that $l_{G}=30\phi^{0.5}$ matches our numerical calculations shown in Figure 10. However, our simulations shown in Fig. 11 suggest a different scaling for the steady state value of $l_{G}$, where $\phi\geq 0.5$ and $\mbox{Pe}\geq 80$. We propose the following phenomenological expression
$l_{G}(\mbox{Pe},\phi)=\zeta(\mbox{Pe})\times\phi^{1.3},$ (15)
where $\zeta(\mbox{Pe}=80)=21$ and $\zeta(\mbox{Pe}=140)=8.5$ are values obtained to fit our data. Using this simple model, we found the following estimations $l_{G}(\mbox{Pe}=80,\phi=0.5)\approx 8.53$, $l_{G}(\mbox{Pe}=140,\phi=0.5)\approx 7.51$, $l_{G}(\mbox{Pe}=80,\phi=0.6)\approx 10.81$, $l_{G}(\mbox{Pe}=140,\phi=0.5)\approx 9.53$, $l_{G}(\mbox{Pe}=80,\phi=0.7)\approx 13.21$, $l_{G}(\mbox{Pe}=140,\phi=0.7)\approx 11.64$, which are in very good agreement with our numerical calculations. For the inner region, we observe a time evolution in the $l_{G}$ value where the average path length increases with particle concentration $\phi$ and reaches a constant value after $t\geq 8t_{r}$ which corresponds to the cluster growth characteristic time reported by Stenhammar et al.stenhammar2014phase .
Figure 11: Average path length for systems in the frontier region (top) and in the inner region (bottom). The values for $l_{G}$ for single-phase systems(top) at $\phi=0.3$ have virtually no change over time and at $\phi=0.4$ increases slowly. In contrast, $l_{G}$ for the separated phase increases rapidly at first and then stabilizes around $t=8t_{r}$.
5 Conclusions
In this work, we have studied the dynamics of ABP using complex networks. We build a time-averaged network and concentrate on the evolution of the topology of the complex network across the MIPS region. We follow the time evolution of the degree distribution, the clustering coefficient, and the average path length, finding different structures of the complex network along the time evolution of the ABP. In the one-phase region, we developed a numerical and theoretical analysis to explain the Gaussian distribution for random graphs. Using a mean free path analysis, we found an analytical expression for the average number of unique encounters of particles in a time window $W$, showing the role of the packing fraction and Péclet number. We reported a transient region, where the degree distribution is hybrid in the sense that it is represented by two components, Gaussian and power laws, namely. Also, in the hybrid region, the average path length and average clustering coefficient do not reach a steady state at times $t=11t_{r}$.
For the two-phase region, we found that the degree distribution follows a double power law distribution with two critical exponents $\gamma_{1}\approx(1.5-3.7)$ and $\gamma_{2}\approx(12-17)$ for $k>6$. We noted that $\gamma_{1}<\gamma_{2}$, being $\gamma_{1}$ more stable over time. Moreover, we analyzed the average clustering coefficient when clustered particles are present. Because of the disk geometry, we analytically showed that this system tries to reach values around $\bar{C}=0.4$ for large packing fractions. The unexpected scaling for the average path length $l_{G}=C(\mbox{Pe})\phi^{1.2}$ is discussed, providing a simple tool to explain our results in the inner region. Our work opens exciting avenues for future research, using complex networks to characterize phase transitions in living organisms, from bacteria and ants to humans. We anticipate our study’s appeal to ecological systems, transport phenomena, multi-agent-based models, and dynamic complex networks. Furthermore, our findings introduce new possibilities for controlling and tracking clusterization in living and artificial active matter. The analytical predictions presented here can serve as a foundation for studying systems with complex interactions, including polarization, alignment, and taxis.

### v2

@@ IV.2 Transient Region: Hybrid phase
We define the transient region as the intermediate phase between the one-phase and two-phase regions (blue in Fig. 2). Since it is an intermediate phase, it is expected to have a degree of distribution with a hybrid behavior. Physically and analyzing our numerical simulations, we note that in this regime, small clusters composed of three disks (triangular arrangement) appear and eventually break because of the persistent ABP motion. In Fig. 5, we observe the time evolution for the degree distribution at times $t=0.45t_{r}$ and $t=30t_{r}$. Interestingly, the initially random graph (Gaussian distribution) evolves to a hybrid distribution with two components (Gaussian and power law). This hybrid behavior tells us that the system is trying to evolve from a random to a scale-free graph, but the values for the packing fraction and Péclet are not large enough. Moreover, at $t=30t_{r}$, the system does not reach the steady state, which is not the case for the one-phase and two-phase regions. The latter tells us that this hybrid phase is more complex to analyze regarding its geometrical and topological properties since the system remains out of equilibrium. In Fig. 2, we colored the data points where we observed this regime in yellow circles.
Figure 5: Time evolution for the degree distribution at the hybrid phase. At the initial time $t=0.45t_{r}$, a Gaussian distribution is observed for the different systems; meanwhile, a hybrid distribution is observed at time $t=30.0t_{r}$. The inset shows the same data plotted in $\log{}-\log{}$ scale.
IV.3 Two-phase region: Double power law distribution
@@
At more significant packing fractions and Péclet numbers, i.e., $\phi\geq 0.5$ and $\mbox{Pe}\geq 80$, the system enters the inner region of Fig. 2, where MIPS is manifested. From the point of view of the ABP, as time evolves, the initially random distribution (liquid-like state) evolves to a state where a gas-like and a solid-like phase cohabit in the system (green-yellow regions in Fig. 2). In terms of the complex network, the degree distribution of the system evolves from a Gaussian distribution at short times ($t=0.45t_{r}$) towards a double power law distribution for larger times ($t\geq 11t_{r}$); see Fig. 6 (a),(b). Notably, when the degree distribution is plotted in the log-log scale, two slopes adequately describe the distribution $P(k)$, see Fig. 6 (c), (d). This double power law distribution has the form
@@
where $K=6$, $k_{c}\in[13-17]$ is the critical degree and $(\gamma_{1},\gamma_{2})$ are the critical exponents. The value $K=6$ is connected with the 2D disk packing problem, where each disk has exactly six nearest neighbors in an ideal cluster and corresponds to the peak of $P(k)$, see Fig. 6 (b). Moreover, the critical values for $k_{c}$ depend on the clustering coefficient and Péclet number. The first slope $\gamma_{1}$ (red dashed line of Fig. 6 (c), (d)) changes from values close to $1.5$ at $\phi=0.5$ towards values close to $3.7$ at $\phi=0.7$. The second slope $\gamma_{2}$ (black dashed lines) varies from values close to $12$ to $17$. In summary, we obtain $\gamma_{1}\approx(1.5-3.7)$ and $\gamma_{2}\approx(12-17)$, and therefore $\gamma_{1}<\gamma_{2}$ in our ABP model. In Table 1, we compare our results with different systems that exhibit a double power law distribution.
@@
The temporal behavior of both curves is shown in Fig. 7 ($\gamma_{1}$) and ($\gamma_{2}$). We note that $\gamma_{1}$ is more stable over time, and the packing fraction dominates its steady value in some inverse relation. On the contrary, $\gamma_{2}$ shows more fluctuations, which can be explained because the clustering formation is unstable for the values of $(\phi,\mbox{Pe})$.
@@
Figure 6: Degree distribution for phase-separated systems. (a) $\log{}-\log{}$ plot shows a clear Gaussian distributions at time $t=0.45t_{r}$. (b) degree distribution at $t=12.0t_{r}$ shows an evolution in the distribution where the degree $k=6$ for each curve is the one with the highest probability. (c) and (d) shows two distributions of figure (B) zoomed in at $\text{Pe},\phi=(80,0.7)$ and $\text{Pe},\phi=(140,0.5)$ respectively, where the red and black dashed lines are the best linear fit for the first and second power law decay respectively for each distribution.
Figure 7: Time evolution for the slopes fitted for the power law distribution (13), in the phase-separated systems.
IV.3.1 Role of the average clustering coefficient
According to our simulations, particles agglomerate when $(\phi,\mbox{Pe})$ cross some critical region, and a natural tendency to form clusters appears. In the language of complex networks, we can elaborate on the connection between this phase separation using the tendency to form clusters from a topological point of view. We encoded the topological information into a dynamical adjacency matrix constructed upon unique encounters in a time interval. Then, we can use the average clustering coefficient presented in Eq. (9) to analyze the global property of such connections to learn how to interpret the inherent hexagonal packing of the solid disks used in our ABP model.
In the upper panel of Fig. 8, we show the temporal behavior of the average clustering coefficient $\bar{C}$ varying the packing fraction ($\phi$) and Péclet number (Pe) and covering different points of the phase diagram illustrated in Fig. 2. In all cases, the natural tendency for a fixed value of Pe is to have higher values of $\bar{C}$ as the packing fraction increases, where $0.1\leq\bar{C}\leq 5.4$. Also, the persistent motion represented by the Péclet number shows that for $\phi\geq 0.5$, an increment of Pe increases does not affect $\bar{C}$ in the long-time behavior, see upper right panel of Fig. 8.
To have an intuition about the higher values of $\bar{C}$, we show in Fig. 9 the average clustering coefficient when particles agglomerate, forming perfect hexagonal clusters. In this ideal cluster, the asymptotic behavior is to reach a value $\bar{C}=0.4$, which can be understood as follows. Let be $\bar{C}_{hex}(N)$ the average clustering coefficient for a hexagonal arrangement with $N$ particles. It is not difficult to show that for $N$ particles, the fractions of particles inside the borders are given by $f_{i}=(1+3N_{r}(N_{r}-1))/(1+3N_{r}(N_{r}+1))$, where $N_{r}=[(1+4/3(N-1))^{1/2}-1]/2$. Moreover, each particle inside the cluster has a clustering coefficient $C_{i}=6/15$. For the particles at the border, we have particles in the corners with clustering coefficient $C_{1}=2/3$, and for the rest of the particles, we have $C_{2}=1/2$. The fraction of particles at the corners is $f_{c}=1/N_{r}$. Based on this, we found the analytical expression:
$\bar{C}_{hex}(N)=\left[f_{c}(N)C_{1}+(1-f_{c}(N))C_{2}\right](1-f_{i}(N))+C_{i}f_{i}(N),$ (14)
which gives us the formula to calculate the average cluster coefficient for an ideal hexagonal cluster. This expression perfectly matches the numerical data points presented in Fig. 9. Most importantly, this tells us that as $N$ increases, the system stabilized around $\bar{C}=0.4$ since $\lim_{N\rightarrow\infty}\bar{C}_{hex}(N)=6/15=0.4$.
IV.3.2 Role of the average path length
@@
The average path length defined in Eq. (10) was measured through all regions, and their values are shown in the lower panel in Fig. 8. We observe the time evolution for the network average path length in the case of a system with a single phase characterized by a Gaussian degree distribution $l_{G}$ has a constant value for different Pe numbers and different packing fractions $\phi$. Meanwhile, on the transient and the two-phase region, the $l_{G}$ value increases slowly in time, reaching a constant value when $t/t_{r}\geq 10$ for the hybrid phase and in $t/t_{r}\geq 8$ in the two-phase region, see Fig 8 (e), (f) respectively. This time scale is similar to the cluster growth characteristic time reported by Stenhammar et al.stenhammar2014phase .
To make progress into understanding this behavior in Fig. 9, we show the average path length in the ideal case of perfect hexagonal clusters with $N$ particles, where we note that a power-law of the form $l_{G}=30\phi^{0.5}$ matches our numerical calculations shown in a magenta dashed line Fig 9. This relation means that while the cluster grows, the network is more connected; therefore, there are more paths connecting particles.
Figure 8: Top: Network average clustering coefficient of identical circles in a hexagonal packing arrangement, using $C_{r}=2^{1/6}\sigma$ ( black circles). $\bar{C}$ approaches asymptotically to $0.4$ (dashed red line), which is the value of the clustering coefficient of the center particle in the seven-particle arrangement (inset). Bottom: Average path length for the same hexagonal packing arrangement as a function of the packing fraction $\phi$ (black circles). We see that $l_{G}$ grows as $30\sqrt{\phi}$ (red dashed line) and could represent an upper bound with respect to $\phi$ for clustered ABP systems.
We propose the following phenomenological expression for the asymptotic values of the average path length in terms of $(\phi,\mbox{Pe})$ plotted in Fig 9,
$l_{G}(\mbox{Pe},\phi)=\alpha+\beta\phi^{\zeta},$ (15)
where $\alpha$, $\beta$ and $\zeta$ are fitting parameters. We present the best-fit results in dashed lines of different colors in Fig. 9 where we observe that for packing fractions $\phi<0.4$, the average path length is approximately constant at different Péclet numbers; this is represented in the fit parameter $\alpha$ which is inversely proportional to Pe. For denser systems $\phi>0.4$, $l_{G}$ increases following a power-law with exponents between $\zeta\approx 2-4$, meanwhile $\beta\approx 20$ for all Pe. This behavior with increasing $\phi$ is a fingerprint of the cluster formation. When a solid phase appears in the steady state of the system, $l_{G}^{\text{solid}}$ contributes significantly to the network connection, suffering a drastic behavior when the phase separation emerges at large Pe.
Figure 9: Top: Network average clustering coefficient $\bar{C}$ as a function of the particle number $N$. Following a crystal shape, $\bar{C}$ reaches an analytical saturation value $\bar{C}\sim 0.4$ as the cluster grows. Bottom: The asymptotic value of the network’s average path length $l_{G}$ as a function of the packing fraction $\phi$. The star symbols represent the simulation’s results for the case of a cluster growing in terms of $\phi$. The dashed magenta line represents the best fit, which follows a power-law behavior. In symbols, we present the asymptotic $l_{G}$ values measured from Fig. 8 in terms of $\phi$ for different Pe. The dashed lines represent the best fits for each case using equation (10).
V Conclusions
In this work, we focused on studying the behavior of Active Brownian Particles (ABP) using complex networks. We examined the changes in the network structure as the packing fraction and Péclet number varied, particularly in the Motility-Induced Phase Separation (MIPS) region. We tracked the evolution of parameters such as the degree distribution, clustering coefficient, and average path length, revealing different network structures as the ABP behavior evolved.
In the single-phase region, we conducted both numerical and theoretical analyses to explain the Gaussian distribution for random graphs. By analyzing the mean free path, we derived an expression for the average number of unique particle encounters within a given time window, highlighting the influence of packing fraction and Péclet number. Surprisingly, even in the single-phase region, we observed a critical packing fraction ($\phi\geq\phi_{c}=\pi/(\pi+2^{13/6})\approx 0.41$) that led to a significant change in network topology. Above this threshold, the mean free path became smaller than the contact radius ($C_{r}$) of the Weeks-Chandler-Andersen potential, indicating the pivotal role of packing fraction in network analysis.
We also identified a transient region where the degree distribution evolved over time, resulting in a hybrid distribution combining Gaussian and power law components. This region corresponded to the formation of small clusters without distinct separation. Our observations did not align with the analytical expression by Redner et al. redner2013structure but matched their numerical predictions and the critical cluster region reported by McDermott et al. mcdermott2023characterizing using machine learning analysis in MIPS. Additionally, the average path length and clustering coefficient stabilized in the hybrid region after $t\geq 11t_{r}$.
In the two-phase region, we found that the degree distribution followed a double power law distribution with two critical exponents, $\gamma_{1}\approx(1.5-3.7)$ and $\gamma_{2}\approx(12-17)$, for $k>6$, representing solid and gas-like phases, respectively. We observed that $\gamma_{1}<\gamma_{2}$, with $\gamma_{1}$ reaching a stable state over time while $\gamma_{2}$ fluctuated around an average value. Furthermore, we analyzed the average clustering coefficient in the presence of clustered particles. Due to the disk geometry of each ABP, we demonstrated that the system tended to approach values around $\bar{C}=0.4$ for large packing fractions. We discussed the behavior of the average path length ($l_{G}$), which increased with the packing fraction and decreased with the Péclet number and provided a heuristic model to describe its behavior across different regions.
Our research paves the way for future studies using complex networks to characterize phase transitions in various living organisms, from bacteria and ants to humans. We believe our findings will be relevant to ecological systems, transport phenomena, multi-agent-based models, and dynamic complex networks. Additionally, our work introduces possibilities for controlling and tracking clusterization in living and artificial active matter and their impact on disease transmission. The analytical predictions presented in this study can form the basis for investigating systems with complex interactions, including polarization, alignment, and taxis.